Is there a stragety to finding a missing numbers on a number line when the missing numbers are the same distance apart??
Example:
2493,3496,4699,5802,605,8008

Answer by checkley77(12844)

(8008-2493)/5=1103 is the missing value to be added.
2493+1103=3596+1103=4699+1103=5802+1103=6905+1103=8008